OpenStudy (anonymous):

simplify the complex rational expressions: show work please √x 1 --- - --- 1 6√x ------------------- √x

OpenStudy (zehanz):

First, write the two fractions in the numerator as one, using\[\frac{ a }{ b }+\frac{ c }{ d }=\frac{ ad+bc }{ bd }\](sounds familiar?) So:\[\frac{ \sqrt{x} }{ 1 }-\frac{ 1 }{ 6\sqrt{x} }=\frac{ 6(\sqrt{x} )^2-1}{ 6\sqrt{x} }=\frac{ 6x-1 }{ 6\sqrt{x} }\]Now, instead of dividing by √x, multiply with the inverse:\[\frac{ 6x-1 }{ 6\sqrt{x} }\cdot \frac{ 1 }{ \sqrt{x} }=\frac{ 6x-1 }{ 6(\sqrt{x})^2 }=\frac{ 6x-1 }{ 6x }\]and you're done!

OpenStudy (anonymous):

the choices are a) x^2- 1/6x b) (√x/1) - (1/6√x) / √x c) 1-1/6 d) 1-1/6x

OpenStudy (abb0t):

Look's like someone answered for you

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

not correct.

OpenStudy (abb0t):

Did you try working it out yourself first? Sometimes answers may not match exactly because they are written in a different form, but essentially mean the same thing.

OpenStudy (anonymous):

go to myalgebra.com and type it in... i have a feeling it is D

OpenStudy (anonymous):

yes i tried but i had no clue what to do and okay

OpenStudy (abb0t):

Peachy, can you show us what you did in your work and we can check where you are stuck on? That way we can guide you in the right direction :)

OpenStudy (anonymous):

we? lol abbot your better than me hehe

OpenStudy (abb0t):

@blondie16 I believe that you are capable of doing this problem too. I have no doubt in my mind.

OpenStudy (anonymous):

lol well than juu hehe

OpenStudy (zehanz):

"not correct" is not correct ;) One could take another step:\[\frac{ 6x-1 }{ 6x }=\frac{ 6x }{ 6x }-\frac{ 1 }{ 6x }=...\]But written as as one fraction also has its charm.